Given that x^2+5x-14 is a factor of x^4-2x^3-37x^2+kx-168, evaluate the sum of the four roots of the equation: x^4-2x^3-37x^2+kx-168=0.
Answer by ikleyn(52832)

.
Given that x^2+5x-14 is a factor of x^4-2x^3-37x^2+kx-168,
evaluate the sum of the four roots of the equation: x^4-2x^3-37x^2+kx-168=0.
~~~~~~~~~~~~~~~~


Independently from any other "given" parts of the problem,


    the sum of the roots of the equation  x^4 - 2x^3 - 37x^2 + kx - 168 = 0 is equal to its coefficient at x^3 
    taken with the opposite sign, i.e. +2.    (Based on the Vieta's theorem)


ANSWER.  The sum of the roots of the equation  x^4 - 2x^3 - 37x^2 + kx - 168 = 0  is  2.


-----------

            Double  and  triple  check  your problem ----- I am  198%  sure that there is a  FATAL  ERROR

            BOTH  in its formulation  AND  in flowing of the thoughts in the head of the person who created it.